OpenStudy (ray_sihota):

Having trouble with this question 27x^2-9x=0 the book is sayin g the answer is x=0 x=1/3. Do not get how they get the answer

OpenStudy (ray_sihota):

9x(3x-1)=0

OpenStudy (anonymous):

\[27 x^2-9 x=0,9 x(3 x-1)=0\] find the two answers by equating each=0

OpenStudy (ray_sihota):

3x-1=0 add 1 to both side 3x=1 x=1/3

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

other 9 x=0,x=0

OpenStudy (ray_sihota):

why is is it 9X0

OpenStudy (ray_sihota):

?

OpenStudy (anonymous):

when x*y=0, then either x=0 or y=0

OpenStudy (anonymous):

I got y=0

OpenStudy (anonymous):

here either 9x=0,x=0 or3 x-1=0 x=1/3

OpenStudy (anonymous):

x=1/3?

OpenStudy (ray_sihota):

got it thanks a lot Surjithayer

OpenStudy (ray_sihota):

so basically factor 9x(3x-1)=0 1st equation 3x-1=0 add 1 to both sides 3x=1 divide by 3X you get 3x=1 then 9x=0 dived by 9 on both sides 0/9= 0

OpenStudy (ray_sihota):

4x^2-x=-9x x(4x-1)=-9x 4x-1=-9x -4x on each side -1=-13x divide by -13 x=1/13

OpenStudy (anonymous):

thanks

OpenStudy (ray_sihota):

but the answer is saying 2 what am I doing wrong is this equation

OpenStudy (ray_sihota):

sorry -2

OpenStudy (anonymous):

yeah its correct

OpenStudy (ray_sihota):

why i am getting 1/13 i do not get -2 like in the back of the book

OpenStudy (anonymous):

\[4 x^2-x=-9 x,4 x^2-x+9 x=0,4x^2+8 x=0,4 x(x+2)=0\] either 4 x=0,x=0/4=0 x+2=0 x=-2